the time flight of a projectile is related to it it's horizontal range by the equation gT2 = 2R the angle of projection is?


1
Expert's answer
2021-07-06T08:23:01-0400

"T=\frac{2v\sin\alpha}{g},"

"R=\frac{2v^2\sin\alpha \cos\alpha}{g},"

"gT^2=2R,"

"\frac{4gv^2\sin^2\alpha}{g^2}=\frac{4v^2\sin\alpha\cos\alpha}{g},"

"\sin\alpha=\cos\alpha,"

"\alpha=45°."